OpenStudy (anonymous):

___2. Find a quadratic function to model the values in the table. Predict the value of y for x = 6.

OpenStudy (anonymous):

A. y = -2x2 + 2x -2; –58 B. y = 2x2 – 2x -2; 60 C. y = 2x2 – 2x -2; 58 D. y = -2x2 + 2x +2; –58

OpenStudy (anonymous):

We need the table.

OpenStudy (anonymous):